OpenStudy (anonymous):

at a bake sale a total of 29 cakes and pies were sold for $283. the price of each cake was $9.50 and the price of each pie was $10. how many cakes and how many pies were sold? at a bake sale a total of 29 cakes and pies were sold for $283. the price of each cake was $9.50 and the price of each pie was $10. how many cakes and how many pies were sold? @Mathematics

OpenStudy (anonymous):

x+y=29 and 9.50x+10y=283 Let's substitute so x=29-y. Plugging that into the 2nd equation equals 9.50(29-y)+10y=283 275.5-9.5y+10y=283. .5y=7.5 y=15 and since 29 in total were sold x=14 So 14 cakes were sold and 15 pies were sold.